# NAMO-LLM: Efficient Navigation Among Movable Obstacles with Large Language Model Guidance

### v1

@@ III-B LLM-guided Sampling Strategy
The sampling-based planner above relies on the distributions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ to generate ${\mathbf{q}}_{\text{new}}$. While these can be uniform, as e.g., in [22], such strategies often lead to slow plan synthesis, particularly as environmental complexity (e.g., the number of movable objects) increases. To accelerate the planning process, we propose non-uniform mass functions $f_{{\mathcal{V}}}$ and $f_{{\mathcal{O}}_{M}}$ that bias sampling toward directions where a feasible solution is more likely to be found. Our approach uses LLMs to identify these promising directions.
@@
where $p_{\text{rand}}\in[0,1)$ is a user-specified parameter and ${\mathcal{V}}_{v}\subseteq{\mathcal{V}}$; ${\mathcal{V}}_{v}$ will be defined later once $f_{{\mathcal{O}}_{M}}$ is defined. According to (1), a node ${\mathbf{q}}\in{\mathcal{V}}_{v}$ is selected to be the node ${\mathbf{q}}_{\text{rand}}$ with probability $p_{\text{rand}}$ while any node ${\mathbf{q}}\in{\mathcal{V}}$ is selected with probability $1-p_{\text{rand}}$. Notice that if $p_{\text{rand}}=0$, then $f_{{\mathcal{V}}}$ models uniform sampling. As $p_{\text{rand}}$ increases, the sampling strategy becomes more biased toward selecting nodes ${\mathbf{q}}\in{\mathcal{V}}_{v}$.
@@
$f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})=\begin{cases}p_{\text{obs}}+\frac{1}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}=o_{\text{LLM}},\\ \frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}\in\mathcal{O}_{M,\text{rand}},\end{cases}$ (2)
where $p_{\text{obs}}\in[0,1)$ is a user-specified parameter and $o_{\text{LLM}}$ is manipulable obstacle, i.e., $o_{\text{LLM}}\in{\mathcal{O}}_{M}$, that is recommended by an LLM to relocate. A discussion on how to prompt an LLM to select $o_{\text{LLM}}$ is deferred to Section III-C. According to (2), the obstacle $o_{\text{LLM}}\in{\mathcal{O}}_{M}$ is selected for relocation with probability $p_{\text{obs}}$ while any obstacle $o_{i}\in{\mathcal{O}}_{M}$ is selected with probability $1-p_{\text{obs}}$. The probability of selecting $o_{i}\notin{\mathcal{O}}_{M}$ is 00. Observe that if $p_{\text{obs}}=0$, then $f_{{\mathcal{O}}_{M}}$ models uniform sampling. As $p_{\text{obs}}$ increases, the sampling strategy becomes more biased toward selecting $o_{\text{LLM}}$. Also, we note that it is possible that the LLM may fail to select a manipulable obstacle or it may generate nonsensical information. In such cases, we set $p_{\text{obs}}=0$ for that iteration so that a manipulable obstacle is selected randomly; see Section III-C.
Next, we define the set ${\mathcal{V}}_{v}$ used for the construction of $f_{{\mathcal{V}}}$ in (1). First, we define a function $v:{\mathcal{V}}\rightarrow\mathbb{N}$ assigning a positive score to each node ${\mathbf{q}}\in{\mathcal{V}}$. This function is initialized as $v({\mathbf{q}}(0))=0$. Second, we show how the score of a newly added node ${\mathbf{q}}_{\text{new}}$ is generated. If ${\mathbf{q}}_{\text{new}}$ was constructed by relocating $o_{\text{LLM}}$, then we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})+1$. Otherwise, we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})$. Informally, $v({\mathbf{q}}_{\text{new}})$ counts how many nodes along the path connecting ${\mathbf{q}}_{\text{new}}$ to ${\mathbf{q}}(0)$ have been constructed based on LLM recommendations. Third, we construct the set ${\mathcal{V}}_{v}$ so that it collects the nodes with the highest score $v$, i.e., ${\mathcal{V}}_{v}=\{{\mathbf{q}}\in{\mathcal{V}}~{}|~{}v({\mathbf{q}})=\max_{{\mathbf{q}}^{\prime}\in{\mathcal{V}}}v({\mathbf{q}}^{\prime})\}$.
Finally, we select a uniform density function for $f_{{\mathcal{C}}_{i}}$. Non-uniform density functions $f_{{\mathcal{C}}_{i}}$ can also be implemented similarly, e.g., by querying an LLM to select configurations $c_{i}^{\prime}$ for which collision-free paths exist; for simplicity, we abstain from this design choice.
@@ III-C Prompt Structure
In this section, we describe how we query a pre-trained LLM to select a manipulable obstacle $o_{\text{LLM}}$. The input prompt to the LLM consists of three parts (see Fig.4): (1) Task description which provides information about the environment’s structure (i.e., the location of all movable and non-movable obstacles and the goal region) and specifies that the robot’s task is to remove obstacles so that the goal region is reachable; (2) Rules that the robot must follow to accomplish its task. Specifically, these rules state that the robot can only remove manipulable obstacles, should minimize the number of obstacles removed and should avoid picking manipulable obstacles that are blocked by other obstacles (see e.g., Case Study VI in Section IV); (3) Output format which defines how the LLM’s response should be formatted for compatibility with Alg. 2. In particular, the LLM is required to generate $K\geq 1$ obstacles in the form $\text{obs}_{i}$ where $i$ is a unique identifier of each obstacle. Among the generated obstacles, one is selected randomly and denoted by $o_{\text{LLM}}$ in (2). Note that the LLM may make content and format errors. A content error occurs when the selected obstacle is not manipulable, and a format error occurs when the LLM’s output does not comply with the required format. If either of these errors occurs, we set $p_{\text{obs}}=0$.
@@ III-D Probabilistic Completeness
In what follows, we show that NAMO-LLM is probabilistically complete. To formally state this result, we first introduce the notion of path clearance [22]. Consider any path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(k),\dots,\tau_{{\mathbf{q}}}(H+1)$ computed by Alg. 1 yielding a plan $\tau$ solving Problem 1. We say that $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$ if any alternative path $\tilde{\tau}_{{\mathbf{q}}}=\tilde{\tau}_{{\mathbf{q}}}(0),\dots,\tilde{\tau}_{{\mathbf{q}}}(k),\dots,\tilde{\tau}_{{\mathbf{q}}}(H+1)$ satisfying the following conditions (i)-(ii) for all $k\in\{1,\dots,H\}$ also corresponds to a feasible plan $\tilde{\tau}$. (a) The transitions from $\tau_{{\mathbf{q}}}(k-1)$ to $\tau_{{\mathbf{q}}}(k)$ and from $\tilde{\tau}_{{\mathbf{q}}}(k-1)$ to $\tilde{\tau}_{{\mathbf{q}}}(k)$ must involve relocating the same manipulable obstacle $o_{i}$, albeit potentially to different configurations. (b) Let $c_{i}^{k}$ and $\tilde{c}_{i}^{k}$ denote the configurations of that obstacle $o_{i}$ in $\tau_{{\mathbf{q}}}(k)$ and $\tilde{\tau}_{{\mathbf{q}}}(k)$, respectively. These configurations must satisfy $\lVert c_{i}^{k}-\tilde{c}_{i}^{k}\rVert\leq\epsilon$ for all $k$, for some $\epsilon>0$. In what follows, we refer to the satisfaction of conditions (i)–(ii) concisely as $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$.
Theorem III.2
Assume that there exists a feasible path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),...,\tau_{{\mathbf{q}}}(H+1)$ with clearance $\epsilon>0$. If $p_{\text{rand}}\in[0,1)$, $p_{\text{obs}}\in[0,1)$, and the motion planner used to check existence of collision-free paths between obstacle configurations in line 2 of Alg. 2 is complete, then the probability that NAMO-LLM will find a feasible path approaches $1$ as $m$ goes to $\infty$.
Proof:
To show this result, we need first to introduce the following notations and definitions. Let ${\mathcal{G}}_{m}=\{{\mathcal{V}}_{m},{\mathcal{E}}_{m}\}$ denote the tree constructed by Algorithm 1 at the end of iteration $m$. Since the plan $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$, there exist alternative feasible plans $\tilde{\tau}_{{\mathbf{q}}}$ such that $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$ for all $k$. Throughout the proof, any reference to a state $\tilde{\tau}_{{\mathbf{q}}}(k)$ assumes this $\epsilon$-closeness to the plan $\tau_{{\mathbf{q}}}$.
@@
Thus we get $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$ completing the proof of part (b). ∎
@@ IV Experiments
In this section, we present extensive numerical experiments to evaluate the efficiency of our method. Specifically, we demonstrate that our approach can effectively solve NAMO problems involving hundreds of movable obstacles and requiring the manipulation of multiple obstacles, significantly outperforming recent search-based methods [18] and sampling-based planners relying on uniform sampling [22] in terms of both runtime and plan quality. Hardware demonstrations of NAMO-LLM on ground robots are provided in [38]. For these experiments, we use the GPT-4o and Gemini-1.5-flash models. All experiments were conducted on a computer with an Intel Core i5 2.4GHz processor and 8 GB of RAM. Software implementation of NAMO-LLM can be found in [37].
@@ IV-A Setting Up Comparative Experiments
NAMO Baselines: We evaluate NAMO-LLM against two baselines: (i) RandomTree, a sampling-based approach from [22] that uses uniform sampling strategies, and (ii) NAMO-SA, a search-based method from [18]. Notably, RandomTree can be recovered from our method by setting $p_{\text{rand}}=p_{\text{obs}}=0$. The approach in [18] selects obstacles for removal by computing the total distance from the robot to the obstacle and from the obstacle to the goal region using the $\text{A}^{*}$ algorithm. This is done for all manipulable objects, and the obstacle with the lowest total distance is selected. The chosen obstacle is then moved to a pre-defined storage room if a collision-free path towards that storage room exists. If no path to the storage room exists, the manipulable obstacle with the next lowest cost is selected; more details can be found in [18]. In our implementation of [18], we model the storage room as an external region adjacent to a specified portion of the workspace boundary. Once an obstacle reaches this boundary segment, it is assumed to be instantaneously transferred to the storage room. On the other hand, our method designs custom drop-off locations for the obstacles during the sampling process. Note that this favors the baseline as our method may sample invalid drop-off locations (e.g., configurations lying in the obstacle space), resulting in sample rejection; see lines 2-2 in Alg. 2. Also in [18], once an obstacle is moved to the storage room, it is never considered for relocation again. This design choice favors that baseline, as the number of obstacles it needs to handle decreases over time—unlike our method. Thus, the maximum horizon $H$ of a plan designed by [18] is equal to the number $n$ of movable obstacles, while in ours it is possible that $H>n$ since our planner is allowed to reconfigure obstacles multiple times. We also note here that [18] includes re-planning strategies to handle unknown obstacles (i.e., obstacles with unknown locations and movability). Since our focus is on known environments, we exclude these capabilities. Finally, to ensure fair comparisons, our implementation of NAMO-LLM ([line 2, Alg. 2]), [18], and [22] share the same RRT planner to check if there exists a collision-free path towards the drop-off location of an obstacle and the storage area.
@@
Robot and Obstacle Configuration Space: We consider a disk-shaped robot with radius $r=0.2$ (Case Studies I-IV) and $r=0.5$ (Case Studies V-VI) and configuration space ${\mathcal{P}}$ determined by the position of the center of the robot. We consider polygonal obstacles with configuration space ${\mathcal{C}}_{i}$ defined by their position and orientation.
Evaluation Metrics: As evaluation metrics, we use (i) the runtime to compute the first feasible plan and (ii) the horizon $H$ of the designed plan. Runtimes and plan horizons $H$ for our method and the sampling-based baseline [22] are averaged over $100$ trials.

### v2

@@ III-B LLM-guided Sampling Strategy
The proposed sampling-based planner relies on the distributions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ to generate ${\mathbf{q}}_{\text{new}}$. While these can be uniform, as in [16], such strategies often result in slow plan synthesis, particularly as environmental complexity (e.g., the number of movable objects) increases. To accelerate planning, we introduce non-uniform mass functions $f_{{\mathcal{V}}}$ and $f_{{\mathcal{O}}_{M}}$ that bias sampling toward directions where feasible solutions are more likely to be found, while retaining a uniform density for $f_{{\mathcal{C}}_{i}}$. To realize this bias, our approach leverages LLMs to identify promising directions.
@@
where $p_{\text{rand}}\in[0,1)$ is a user-specified parameter and ${\mathcal{V}}_{v}\subseteq{\mathcal{V}}$; ${\mathcal{V}}_{v}$ will be defined later once $f_{{\mathcal{O}}_{M}}$ is defined. Notice that if $p_{\text{rand}}=0$, then $f_{{\mathcal{V}}}$ models uniform sampling while as $p_{\text{rand}}$ increases, the sampling strategy becomes more biased toward selecting nodes ${\mathbf{q}}\in{\mathcal{V}}_{v}$.
@@
$f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})=\begin{cases}p_{\text{obs}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|}},&\text{if}~o_{i}=o_{\text{LLM}},\\ \frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~o_{i}\in\mathcal{O}_{M,\text{rand}},\end{cases}$ (2)
where, $p_{\text{obs}}\in[0,1)$ is a user-specified parameter, and $o_{\text{LLM}}\in{\mathcal{O}}_{M}$ is a manipulable obstacle recommended by an LLM for relocation. How to prompt the LLM to select $o_{\text{LLM}}$ is discussed in Section III-C. According to (2), the probability of selecting $o_{i}\notin{\mathcal{O}}_{M}$ is 0. If $p_{\text{obs}}=0$, $f_{{\mathcal{O}}_{M}}$ reduces to uniform sampling; as $p_{\text{obs}}$ increases, sampling increasingly favors $o_{\text{LLM}}$; see Fig. 3. If the LLM fails to select a manipulable obstacle or outputs invalid information, we set $p_{\text{obs}}=0$ for that iteration so a manipulable obstacle is chosen randomly; see Section III-C.
Next, we define the set ${\mathcal{V}}_{v}$ used to construct $f_{{\mathcal{V}}}$ in (1). First, we define a function $v:{\mathcal{V}}\rightarrow\mathbb{N}$ assigning a positive score to each tree node $\mathbf{q}$}. This function is initialized as $v({\mathbf{q}}(0))=0$. Second, we show how the score of a newly added node ${\mathbf{q}}_{\text{new}}$ is generated; see Fig. 2. If ${\mathbf{q}}_{\text{new}}$ was constructed by relocating $o_{\text{LLM}}$, then we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})+1$. Otherwise, we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})$. Informally, $v({\mathbf{q}}_{\text{new}})$ counts how many nodes along the path connecting ${\mathbf{q}}_{\text{new}}$ to ${\mathbf{q}}(0)$ have been constructed based on LLM recommendations. Third, we construct ${\mathcal{V}}_{v}$ so that it collects the nodes with the highest score, i.e., ${\mathcal{V}}_{v}=\{{\mathbf{q}}\in{\mathcal{V}}~|~v({\mathbf{q}})=\max_{{\mathbf{q}}^{\prime}\in{\mathcal{V}}}v({\mathbf{q}}^{\prime})\}$; see Fig. 2.
@@ III-C Prompt Structure
In this section, we describe how we prompt a pre-trained LLM to select a manipulable obstacle $o_{\text{LLM}}$ in a single query. The input prompt has three parts (see Fig.4): (1) Task description, which specifies the environment (locations of movable/non-movable obstacles and the goal region) and the robot’s task of removing obstacles to reach the goal; (2) Rules, which state that the robot may only remove manipulable obstacles, should minimize removals, and should avoid selecting blocked manipulable obstacles (see Case StudyVI in Section IV); and (3) Output format, which defines how the LLM’s response must be structured for compatibility with Alg.2. In particular, the LLM is asked to generate $K\geq 1$ obstacles (in a single query) in the form $\text{obs}i$, where $i$ uniquely identifies an obstacle. One of the generated obstacles is then selected at random and denoted $o_{\text{LLM}}$ in (2). If $o_{\text{LLM}}$ is not manipulable (content error) or the response does not follow the required structure (format error), we discard the LLM output and set $p_{\text{obs}}=0$ for that iteration.
Figure 4: Input prompt for GPT-4o and $K=1$ in Case Study V (Sec. IV).
@@ III-D Probabilistic Completeness
Next, we show that NAMO-LLM is probabilistically complete. To formally state this result, we first introduce the notion of path clearance [16]. Consider any path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(k),\dots,\tau_{{\mathbf{q}}}(H+1)$ computed by Alg. 1 yielding a plan $\tau$ solving Problem 1. We say that $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$ if any alternative path $\tilde{\tau}_{{\mathbf{q}}}=\tilde{\tau}_{{\mathbf{q}}}(0),\dots,\tilde{\tau}_{{\mathbf{q}}}(k),\dots,\tilde{\tau}_{{\mathbf{q}}}(H+1)$ satisfying the following conditions (i)-(ii) for all $k\in\{1,\dots,H\}$ also corresponds to a feasible plan $\tilde{\tau}$. (a) The transitions from $\tau_{{\mathbf{q}}}(k-1)$ to $\tau_{{\mathbf{q}}}(k)$ and from $\tilde{\tau}_{{\mathbf{q}}}(k-1)$ to $\tilde{\tau}_{{\mathbf{q}}}(k)$ must involve relocating the same manipulable obstacle $o_{i}$, albeit potentially to different configurations. (b) Let ${\mathbf{c}}_{i}^{k}$ and $\tilde{{\mathbf{c}}}_{i}^{k}$ denote the configurations of that obstacle $o_{i}$ in $\tau_{{\mathbf{q}}}(k)$ and $\tilde{\tau}_{{\mathbf{q}}}(k)$, respectively. These configurations must satisfy $\lVert{\mathbf{c}}_{i}^{k}-\tilde{{\mathbf{c}}}_{i}^{k}\rVert\leq\epsilon$ for all $k$, for some $\epsilon>0$. In what follows, we refer to satisfaction of conditions (i)–(ii) concisely as $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$.
Theorem III.1
Assume that there exists a feasible path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),...,\tau_{{\mathbf{q}}}(H+1)$ with clearance $\epsilon>0$. If $p_{\text{rand}}\in[0,1)$, $p_{\text{obs}}\in[0,1)$, and the motion planner used to check existence of collision-free paths in [line 2, Alg. 2] is complete, then the probability that NAMO-LLM will find a feasible path approaches $1$ as $m$ goes to $\infty$.
Proof: To show this result, we need first to introduce the following notations and definitions. Let ${\mathcal{G}}_{m}=\{{\mathcal{V}}_{m},{\mathcal{E}}_{m}\}$ denote the tree constructed by Algorithm 1 at the end of iteration $m$. Since the plan $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$, there exist alternative feasible plans $\tilde{\tau}_{{\mathbf{q}}}$ such that $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$ for all $k$. Throughout the proof, any reference to a state $\tilde{\tau}_{{\mathbf{q}}}(k)$ assumes this $\epsilon$-closeness to the plan $\tau_{{\mathbf{q}}}$.
@@
Thus we get $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$ completing the proof of part (b).
@@ IV Experiments
In this section, we demonstrate that NAMO-LLM outperforms search-based methods [12], sampling-based planners [16], and LLM-based planners in terms of runtimes, plan quality, and task success rates. Our evaluation is performed across different LLMs and environmental setups. Hardware demonstrations of our method on ground robots are also provided; see also the supplemental material. All experiments were conducted on a computer with an Intel Core i5 2.4GHz processor and 8 GB of RAM. Software implementation of NAMO-LLM can be found in [29].
@@ IV-A Setting Up Comparative Experiments
NAMO Baselines: We consider three baselines: (i) RandomTree, a sampling-based approach from [16] that employs uniform sampling; (ii) NAMO-SA, a search-based method from [12]; and (iii) an LLM-based planner that relies exclusively on LLMs. Among these, only (iii) incorporates LLMs. Comparisons with (i)–(ii) are reported in Sections IV-BIV-C, and comparisons with (iii) are presented in Section IV-D.
RandomTree is recovered from our method by setting $p_{\text{rand}}=p_{\text{obs}}=0$, while the LLM-based planner corresponds to NAMO-LLM with $p_{\text{rand}}=p_{\text{obs}}=1$ and $K=1$. Unlike our method and RandomTree, the LLM-based planner lacks completeness guarantees (see Theorem III.1). RandomTree and NAMO-LLM run until either a feasible solution is found or a maximum of $m_{\text{max}}=30,000$ iterations is reached. In the LLM-based planner, the LLM is queried at every iteration to select an obstacle, which is placed at a random location. To ensure financially sustainable but fair comparisons, we cap the total number of LLM queries to $3H_{\text{min}}$, where $H_{\text{min}}$ is the minimum number of obstacles that need to be relocated to clear a path to the goal area. At each iteration, the LLM receives the same prompt as NAMO-LLM and outputs an obstacle. If the output has no format/content errors, the obstacle is placed randomly; otherwise, the counter is incremented and the LLM is queried again.
@@
The approach in [12] selects obstacles for removal by computing the total distance from the robot to the obstacle and from the obstacle to the goal region using $\text{A}^{*}$. The obstacle with the lowest cost is chosen and moved to a pre-defined storage room if a collision-free path exists; otherwise, the next lowest-cost obstacle is considered. In our implementation, the storage room is modeled as an external region adjacent to the workspace boundary, and once an obstacle reaches this boundary it is assumed to be instantaneously transferred. In contrast, our method designs custom drop-off locations during sampling. This favors the baseline, since our method may sample invalid drop-off locations leading to rejections; see lines 22 in Alg. 2. Moreover, in [12], once an obstacle is moved to the storage room it is never relocated, further favoring the baseline as the number $n$ of movable obstacles decreases over time—unlike in our method. Thus, the maximum horizon $H$ of a plan in [12] equals $n$, whereas in ours it is possible that $H>n$ since obstacles may be reconfigured multiple times. Re-planning strategies in [12] to handle unknown obstacles are neglected in our implementation as we focus on known environments. Finally, to ensure fair comparisons, our implementation of NAMO-LLM ([line 2, Alg. 2]), [12], and [16] share the same RRT planner to check if there exists a collision-free path towards the drop-off location of an obstacle and the storage area.
@@
Robot and Obstacle Configuration Space: We consider a disk-shaped robot of radius $r$ ($r=0.2$ in Cases I–IV, $r=0.5$ in Cases V–VI) with configuration space $\mathcal{P}$} given by the robot’s center position. Obstacles are polygonal, each with configuration space ${\mathcal{C}}_{i}$ defined by its position and orientation.
Evaluation Metrics: As evaluation metrics we use (a) the runtime to compute the first feasible plan; (b) the horizon $H$ of the designed plan; and (c) task success rates. Task success rates, runtimes, and plan horizons $H$ for our method, the LLM-based planner, and [16] are averaged over $100$ runs. Task success rates are defined as the percentage of runs that result in a feasible manipulation plan. Note that NAMO-SA’s success rate is always $100\%$, since it has no randomized component and, in the worst case, moves all objects sequentially to a storage room as discussed earlier. RandomTree and NAMO-LLM will also have a $100\%$ success rate, for large enough $m_{\text{max}}$, due to their completeness properties. Our reported times include all computational components of Alg. 1–2 including e.g., the runtime for LLM inference.
Figure 5: Initial configurations of the obstacles (red polygons) and robot for Case Studies I-IV. The green path connects the initial robot state to the first obstacle to be repositioned for clearing a path to the goal area ${\mathcal{P}}_{g}$.
IV-B Comparisons with Search- and Sampling-Based Baselines